# Towards Unified and Lossless Latent Space for 3D Molecular Latent Diffusion Modeling

### v3

@@ Abstract
3D molecule generation is crucial for drug discovery and material science, requiring models to process complex multi-modalities, including atom types, chemical bonds, and 3D coordinates. A key challenge is integrating these modalities of different shapes while maintaining SE(3) equivariance for 3D coordinates. To achieve this, existing approaches typically maintain separate latent spaces for invariant and equivariant modalities, reducing efficiency in both training and sampling. In this work, we propose Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling (UAE-3D), a multi-modal VAE that compresses 3D molecules into latent sequences from a unified latent space, while maintaining near-zero reconstruction error. This unified latent space eliminates the complexities of handling multi-modality and equivariance when performing latent diffusion modeling. We demonstrate this by employing the Diffusion Transformer–a general-purpose diffusion model without any molecular inductive bias–for latent generation. Extensive experiments on GEOM-Drugs and QM9 datasets demonstrate that our method significantly establishes new benchmarks in both de novo and conditional 3D molecule generation, achieving leading efficiency and quality. On GEOM-Drugs, it reduces FCD by 72.6% over the previous best result, while achieving over 70% relative average improvements in geometric fidelity.
@@ 1 Introduction
Figure 1: (a) A 3D molecule has multi-modal features. (b) Prior methods use separate latent spaces for equivariant (3D) and invariant (2D) modalities, inducing unnecessary complexity for the model architecture. (c) UAE-3D reduces this complexity by establishing a unified and lossless latent space that integrates all molecular modalities.
@@
Despite its simplicity, UAE-3D can compress 3D molecules into token sequences in a unified latent space, eliminating the complexities of handling multi-modalities and 3D equivariance in latent diffusion modeling. Figure 2(a;b) show that UAE-3D reaches lossless reconstruction of atom and bond types, with a near-zero coordinate RMSD. To further demonstrate its effectiveness, we employ the Diffusion Transformer (DiT) [23], a general-purpose diffusion backbone without any molecular inductive bias, to model UAE-3D’s latents. We show that DiT can successfully generate stable and valid 3D molecules, with significantly improved training (by 2.7 times) and sampling (by 7.3 times) efficiency (cf. Figure 2(c;d)). We refer to this LDM pipeline as UDM-3D: Unified Latent Diffusion Modeling for 3D Molecule Generation.
@@ 2 Background: 3D Molecular Latent Diffusion Models
[…]
@@ 3.1 Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling
$\mathbf{Q}_{ij}=[\mathbf{H}^{n}_{i};\mathbf{H}^{e}_{ij}]\mathbf{W}^{q}\in\mathbb{R}^{d},$ (5) $[\mathbf{K}_{ij};\mathbf{V}_{ij}]=[\mathbf{H}^{n}_{j};\mathbf{H}^{e}_{ij}]\mathbf{W}^{kv}\in\mathbb{R}^{2d},$ (6)
@@
SE(3)-Equivariant Augmentations. To enforce SE(3)-equivariance in UAE-3D, we apply random transformations $\mathbf{R}\in\text{SE}(3)$ to input coordinates $\mathbf{X}$ during training. Each $\mathbf{R}$ combines a rotation from $\text{SO}(3)$ and a random translation from $\mathcal{N}(\mathbf{0},0.01\mathbf{I}_{3})$. Crucially, the reconstruction loss $\mathcal{L}_{\text{recon}}$ (cf. Equation (12)) remains unchanged but operates on transformed inputs $\mathbf{R}\circ\mathbf{G}=\langle\mathbf{F},\mathbf{E},\mathbf{R}(\mathbf{X})\rangle$. This process can be informally conceptualized as a reconstruction loss of $\|\mathcal{D}(\mathcal{E}(\mathbf{R}\circ\mathbf{G}))-\mathbf{R}\circ\mathbf{G}\|$, where the autoencoder $\mathcal{D}(\mathcal{E}(\cdot))$ learns to preserve geometric consistency under SE(3) transformations.
@@ 3.2 Unified Latent Diffusion Modeling for 3D Molecule Generation
Diffusion Transformer (DiT). Given the unified latent space provided by our UAE-3D model, we adopt the DiT [23] as the backbone diffusion model $\boldsymbol{\epsilon}_{\theta}$ for 3D molecular latent generation. Originally developed for image LDM, DiT has demonstrated strong performance in modeling latent sequences. Compared to a standard Transformer [22], DiT replaces the layernorm [28] by adaptive layernorm (adaLN) [29], where the scale $\mathbf{y}\in\mathbb{R}^{d}$ and shift $\mathbf{b}\in\mathbb{R}^{d}$ parameters are conditioned on the diffusion timestep $t\in(0,1]$. Specifically, $[\mathbf{y};\mathbf{b}]=\text{MLP}_{\text{adaLN}}(\text{Embed}_{t}(t))$, where $\text{Embed}_{t}(\cdot)$ is a shared linear layer and $\text{MLP}_{\text{adaLN}}$ is a module specific to each adaLN instance. These generated shift and scale parameters act as “soft gates”, enabling timestep-dependent activation of DiT’s hidden representations and facilitating more effective, scale-adaptive denoising.
@@
$[\mathbf{y};\mathbf{b}]=\text{MLP}_{\text{adaLN}}(\text{Embed}_{t}(t)+\text{Embed}_{c}(\mathbf{c})).$ (14)
@@
${}_{\text{3D}}$$\downarrow$ AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
@@
[…]
@@ 4.1 Experimental Setup
Baselines. For de novo 3D molecule generation, we compare UDM-3D with CDGS [32], JODO [6], MiDi [9], G-SchNet [12], G-SphereNet [33], EDM [5], MDM [34], GeoLDM [17], and ADiT [35]. For conditional generation, we additionally use baselines of EEGSDE [8] and GeoBFN [36]. We do not report MDM’s performance on GEOM-DRUGS because we are unable to reproduce reasonable results using their official code. Similarly, ADiT’s performance on GEOM-DRUGS is omitted, because it is not available in their paper, code, and released checkpoint.
@@ 4.2 De Novo 3D Molecule Generation
Experimental Setting. Generating a 3D molecule involves generating the structural validity (atom/bond features) and accurate spatial arrangements (3D coordinates). Therefore, we adapt the comprehensive metrics from [6, 5]. We also report all the metrics for molecules from training set, serving as an approximate upper bound for performance. Our evaluation metrics can be divided into two groups: (1) 2D Metrics: Atom stability, validity & completeness (V&C), validity & uniqueness (V&U), validity & uniqueness & novelty (V&U&N), similarity to nearest neighbor (SNN), fragment similarity (Frag), scaffold similarity (Scaf), and Fréchet ChemNet Distance (FCD) [3]; (2) 3D Metrics: Atom stability, FCD, maximum mean discrepancy (MMD) [37] for the distributions of bond lengths, bond angles, and dihedral angles. More experimental details are in Appendix D.1.
Table 2: Performance of de novo 3D molecule generation on QM9. * indicates results reproduced using official source codes, while other baseline results are taken from [6]. ADiT’s relatively poor performance stems from differences in evaluation protocols -— their generated molecules exhibit significant stability issues. We provide a detailed analysis in Appendix D
@@
${}_{\text{3D}}$$\downarrow$ AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
@@
ADiT* 2.884 0.211 0.000 9.98E-01 3.38E-02 1.46E-03
[…]
@@
Breakthrough in Geometric Accuracy. Our unified latent space enables highly accurate 3D geometric modeling, a key factor in realistic molecule generation. UDM-3D significantly outperforms baselines in distributional distances of bond lengths and angles, reducing errors by an order of magnitude. These improvements hold even for complex, drug-like molecules in GEOM-Drugs—where UDM-3D achieves a 25x lower bond length error than GeoLDM (9.89E-03 v.s. 3.91E-01). This leap in accuracy comes from UDM-3D’s ability to jointly optimize chemical and geometric constraints in a unified latent space, maintaining consistency across molecular modalities.
@@ 4.3 Conditional 3D Molecule Generation
Experimental Settings. We evaluate conditional generation of molecules with target quantum properties using the protocol from [5, 17]. Specifically, our target properties include $C_{v}$, $\mu$, $\alpha$, $\epsilon_{\text{HOMO}}$, $\epsilon_{\text{LUMO}}$, $\Delta\epsilon$, and we evaluate the Mean Absolute Error (MAE) between target and predicted properties. More details on the properties and settings are provided in Appendix D.2.
@@
[…]
@@ 4.4 Training & Sampling Efficiency
Table 4: Comparing DiT with Transformer for latent diffusion modeling. Models have the same depth and hidden size. A.S. $\rightarrow$ atom stability.
Metric DiT Transformer
A.S.${}_{\text{3D}}$ 0.993 0.983
A.S.${}_{\text{2D}}$ 0.999 0.997
V&C 0.983 0.938
V&U 0.973 0.922
V&U&N 0.950 0.922
@@
${}_{\text{3D}}$$\downarrow$ AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
@@ 4.5 Analysis and Ablation Studies
Ablating the DiT Backbone. We compare DiT with a vanilla Transformer for de novo 3D molecule generation on the QM9 dataset (cf. Table 4). The results show that DiT outperforms the vanilla Transformer across all metrics, which we attribute to its adaptive LayerNorm layers. These layers enable DiT to effectively handle data with varying noise scales, improving diffusion performance.
@@
Train with RMSD (×$\times 10^{-4}$Å)
@@ 6 Conclusion and Future Works
In this paper, we propose UAE-3D to compress the multi-modal features of 3D molecules into a unified latent space, and demonstrate the effectiveness of latent diffusion modeling on this space by introducing UDM-3D. By integrating atom types, chemical bonds, and 3D coordinates into a single latent space, our model effectively addresses the inherent challenges of multi-modality and SE(3) equivariance for 3D molecule generation. Extensive experiments on GEOM-Drugs and QM9 confirm leading performances in both de novo and conditional 3D molecule generation, setting new benchmarks for both quality and efficiency. UAE-3D’s success highlights the benefits of building a unified latent space for molecular design. Moving forward, we will transfer this unified latent space to new modalities, including proteins and RNAs, and extend it to broader molecular design tasks, including structure-based drug design.
Appendix A Limitations and More Future Works
@@
Future Works: Multi-modal Molecule Language Models. While UAE-3D’s latents are near-lossless compressions of the original molecule, it inspires the exploration of jointly modeling UAE-3D’s latents with text sequences for text-guided 3D molecule generation, following [42, 43, 44]
Appendix B More Experimental Results
B.1 More Ablation Studies and Analysis
@@
4 0.914 0.325 0.580 0.566 0.566 3.27E-01 2.20E-01 4.19E-03
8 0.987 0.882 0.942 0.923 0.923 1.29E-01 1.77E-02 8.14E-04
16 0.999 0.988 0.983 0.973 0.950 7.04E-02 9.84E-03 3.47E-04
32 0.986 0.867 0.943 0.918 0.918 9.98E-02 1.42E-02 7.77E-04
@@
$\lambda$ / ${}_{\text{3D}}$$\downarrow$ AtomStable MolStable Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
@@
Influenece of Latent Dimensions. To investigate the impact of latent dimensionality on 3D molecule generation, we conduct an ablation study on the QM9 dataset. Table 7 summarizes the performance metrics across different latent space dimensions (4, 8, 16, and 32). The results indicate that a too-low latent dimension (e.g., 4) fails to capture the necessary information, leading to inferior performance. Increasing the dimension to 16 results in substantial performance improvements, while further increasing to 32 does not yield significant additional gains. This demonstrates that a latent dimension of 16 provides an optimal trade-off between model capacity and reconstruction fidelity.
@@
B.2 More Visualizations of Generated Molecules

### v4

@@ Abstract
3D molecule generation is crucial for drug discovery and material science, requiring models to process complex multi-modalities, including atom types, chemical bonds, and 3D coordinates. A key challenge is integrating these modalities of different shapes while maintaining SE(3) equivariance for 3D coordinates. To achieve this, existing approaches typically maintain separate latent spaces for invariant and equivariant modalities, reducing efficiency in both training and sampling. In this work, we propose Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling (UAE-3D), a multi-modal VAE that compresses 3D molecules into latent sequences from a unified latent space, while maintaining near-zero reconstruction error. This unified latent space eliminates the complexities of handling multi-modality and equivariance when performing latent diffusion modeling. We demonstrate this by employing the Diffusion Transformer–a general-purpose diffusion model without any molecular inductive bias–for latent generation. Extensive experiments on GEOM-Drugs and QM9 datasets demonstrate that our method significantly establishes new benchmarks in both de novo and conditional 3D molecule generation, achieving leading efficiency and quality. On GEOM-Drugs, it reduces FCD by 72.6% over the previous best result, while achieving over 70% relative average improvements in geometric fidelity. Our code is released at https://github.com/lyc0930/UAE-3D/.
@@ 1 Introduction
Figure 1: (a) A 3D molecule has multi-modal features. (b) Prior methods use separate latent spaces for equivariant (3D) and invariant (2D) modalities, inducing unnecessary complexity for the model architecture. (c) UAE-3D reduces this complexity by establishing a unified and near-lossless latent space that integrates all molecular modalities.
@@
Despite its simplicity, UAE-3D can compress 3D molecules into token sequences in a unified latent space, eliminating the complexities of handling multi-modalities and 3D equivariance in latent diffusion modeling. Figure 2(a;b) show that UAE-3D reaches near-lossless reconstruction of atom and bond types, with a near-zero coordinate RMSD. To further demonstrate its effectiveness, we employ the Diffusion Transformer (DiT) [23], a general-purpose diffusion backbone without any molecular inductive bias, to model UAE-3D’s latents. We show that DiT can successfully generate stable and valid 3D molecules, with significantly improved training (by 2.7 times) and sampling (by 7.3 times) efficiency (cf. Figure 2(c;d)). We refer to this LDM pipeline as UDM-3D: Unified Latent Diffusion Modeling for 3D Molecule Generation.
@@ 2 Background: 3D Molecular Latent Diffusion Models
Separated Latent Spaces. The VAE latents in previous works [18, 17] include two parts: $\mathbf{Z}=[\bar{\mathbf{Z}};\vec{\mathbf{Z}}]$, where $\vec{\mathbf{Z}}$ is for equivariant features and $\bar{\mathbf{Z}}$ is for 3D invariant features. They also rely on diffusion models with baked-in 3D equivariance. In contrast, as Section 3 shows, our method uses a unified latent space, allowing a general-purpose diffusion model–without any geometric or molecular inductive bias–to achieve strong performance.
@@ 3.1 Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling
$\mathbf{Q}_{ij}=[\mathbf{H}^{n}_{i};\mathbf{H}^{e}_{ij}]\mathbf{W}^{q}\in\mathbb{R}^{d},$ (5) $\mathbf{K}_{ij};\mathbf{V}_{ij}]=[\mathbf{H}^{n}_{j};\mathbf{H}^{e}_{ij}]\mathbf{W}^{kv}\in\mathbb{R}^{2d},$ (6)
@@
SE(3)-Equivariant Augmentations. To enforce SE(3)-equivariance in UAE-3D, we apply random transformations $\mathbf{R}\in\text{SE}(3)$ to input coordinates $\mathbf{X}$ during training [28, 20]. Each $\mathbf{R}$ combines a rotation from $\text{SO}(3)$ and a random translation from $\mathcal{N}(\mathbf{0},0.01\mathbf{I}_{3})$. Crucially, the reconstruction loss $\mathcal{L}_{\text{recon}}$ (cf. Equation (12)) remains unchanged but operates on transformed inputs $\mathbf{R}\circ\mathbf{G}=\langle\mathbf{F},\mathbf{E},\mathbf{R}(\mathbf{X})\rangle$. This process can be informally conceptualized as a reconstruction loss of $\|\mathcal{D}(\mathcal{E}(\mathbf{R}\circ\mathbf{G}))-\mathbf{R}\circ\mathbf{G}\|$, where the autoencoder $\mathcal{D}(\mathcal{E}(\cdot))$ learns to preserve geometric consistency under SE(3) transformations.
@@ 3.2 Unified Latent Diffusion Modeling for 3D Molecule Generation
Diffusion Transformer (DiT). Given the unified latent space provided by our UAE-3D model, we adopt the DiT [23] as the backbone diffusion model $\boldsymbol{\epsilon}_{\theta}$ for 3D molecular latent generation. Originally developed for image LDM, DiT has demonstrated strong performance in modeling latent sequences. Compared to a standard Transformer [22], DiT replaces the layernorm [29] by adaptive layernorm (adaLN) [30], where the scale $\mathbf{y}\in\mathbb{R}^{d}$ and shift $\mathbf{b}\in\mathbb{R}^{d}$ parameters are conditioned on the diffusion timestep $t\in(0,1]$. Specifically, $\mathbf{y};\mathbf{b}]=\text{MLP}_{\text{adaLN}}(\text{Embed}_{t}(t))$, where $\text{Embed}_{t}(\cdot)$ is a shared linear layer and $\text{MLP}_{\text{adaLN}}$ is a module specific to each adaLN instance. These generated shift and scale parameters act as “soft gates”, enabling timestep-dependent activation of DiT’s hidden representations and facilitating more effective, scale-adaptive denoising.
@@
$\mathbf{y};\mathbf{b}]=\text{MLP}_{\text{adaLN}}(\text{Embed}_{t}(t)+\text{Embed}_{c}(\mathbf{c})).$ (14)
@@
[…]
@@
UDM-3D, ours ${\textbf{17.36}}^{\color[rgb]{1,.5,0}-13.2\%}$ 0.852 0.014 $\text{{{9.89E-03}}}^{\color[rgb]{1,.5,0}-88.4\%}$ $\text{{{5.11E-03}}}^{\color[rgb]{1,.5,0}-55.6\%}$ $\text{{{1.74E-04}}}^{\color[rgb]{1,.5,0}-74.0\%}$
@@ 4.1 Experimental Setup
Baselines. For de novo 3D molecule generation, we compare UDM-3D with CDGS [33], JODO [6], MiDi [9], G-SchNet [12], G-SphereNet [34], EDM [5], MDM [35], GeoLDM [17], EQGAT-diff [36], SemlaFlow[37] and ADiT [38]. For conditional generation, we additionally use baselines of EEGSDE [8] and GeoBFN [39]. We do not report MDM’s performance on GEOM-DRUGS because we are unable to reproduce reasonable results using their official code. Similarly, ADiT’s performance on GEOM-DRUGS is omitted, because it is not available in their paper, code, and released checkpoint.
@@ 4.2 De Novo 3D Molecule Generation
Experimental Setting. Generating a 3D molecule involves generating the structural validity (atom/bond features) and accurate spatial arrangements (3D coordinates). Therefore, we adapt the comprehensive metrics from [6, 5]. We also report all the metrics for molecules from training set, serving as an approximate upper bound for performance. Our evaluation metrics can be divided into two groups: (1) 2D Metrics: Atom stability, validity & completeness (V&C), validity & uniqueness (V&U), validity & uniqueness & novelty (V&U&N), similarity to nearest neighbor (SNN), fragment similarity (Frag), scaffold similarity (Scaf), and Fréchet ChemNet Distance (FCD) [3]; (2) 3D Metrics: Atom stability, FCD, maximum mean discrepancy (MMD) [40] for the distributions of bond lengths, bond angles, and dihedral angles. More experimental details are in Appendix C.1.
Table 2: Performance of de novo 3D molecule generation on QM9. * indicates results reproduced using official source codes, while other baseline results are taken from [6]. Some of ADiT’s evaluation metrics are omitted due to differences in evaluation protocols. We discuss these protocol differences and their implications in Appendix C.
@@
[…]
@@
SemlaFlow 1.127 0.971 0.787 - - -
ADiT* 2.884 0.211 - 9.98E-01 3.38E-02 1.46E-03
UDM-3D, ours ${\textbf{0.881}}^{\color[rgb]{1,.5,0}-0.45\%}$ 0.993 0.935 $\text{{{7.04E-02}}}^{\color[rgb]{1,.5,0}-45.8\%}$ $\text{{{9.84E-03}}}^{\color[rgb]{1,.5,0}-1.6\%}$ $\text{{{3.47E-04}}}^{\color[rgb]{1,.5,0}-44.8\%}$
@@
Leading Geometric Accuracy. Our unified latent space enables highly accurate 3D geometric modeling, a key factor in realistic molecule generation. UDM-3D significantly outperforms baselines in distributional distances of bond lengths and angles, reducing errors by an order of magnitude. These improvements hold even for complex, drug-like molecules in GEOM-Drugs—where UDM-3D achieves a 25x lower bond length error than GeoLDM (9.89E-03 v.s. 3.91E-01). This leap in accuracy comes from UDM-3D’s ability to jointly optimize chemical and geometric constraints in a unified latent space, maintaining consistency across molecular modalities.
@@ 4.3 Conditional 3D Molecule Generation
Experimental Settings. We evaluate conditional generation of molecules with target quantum properties using the protocol from [5, 17]. Specifically, our target properties include $C_{v}$, $\mu$, $\alpha$, $\epsilon_{\text{HOMO}}$, $\epsilon_{\text{LUMO}}$, $\Delta\epsilon$, and we evaluate the Mean Absolute Error (MAE) between target and predicted properties. More details on the properties and settings are provided in Appendix C.2.
@@
UDM-3D, ours ${\textbf{0.603}}^{\color[rgb]{1,.5,0}-3.98\%}$ $1.54^{\color[rgb]{.5,.5,.5}+8.4\%}$ ${\textbf{0.553}}^{\color[rgb]{1,.5,0}-4.82\%}$ ${\textbf{216}}^{\color[rgb]{1,.5,0}-4.42\%}$ ${\textbf{\phantom{0}247}}^{\color[rgb]{1,.5,0}-3.5\%} 0$${\textbf{\phantom{0}313}}^{\color[rgb]{1,.5,0}-6.57\%}$
@@ 4.4 Training & Sampling Efficiency
Table 4: Comparing DiT with Transformer and PerceiverIO for latent diffusion modeling. Models have the same depth and hidden size. A.S. $\rightarrow$ atom stability.
Metric DiT Transformer PerceiverIO
[…]
V&C 0.983 0.938 0.933
V&U 0.973 0.922 0.931
V&U&N 0.950 0.922 0.931
@@
[…]
@@ 4.5 Analysis and Ablation Studies
Ablating the DiT Backbone. We compare DiT with alternative transformer-based architectures for de novo 3D molecule generation on the QM9 dataset (cf. Table 4). Specifically, we evaluate against a vanilla Transformer and PerceiverIO[41], a modern architecture designed for structured inputs and outputs. The results show that DiT consistently outperforms both baselines across all metrics, which we attribute to its adaptive LayerNorm layers. These layers enable DiT to effectively handle data with varying noise scales, thereby improving diffusion performance. Moreover, when paired with alternative diffusion neural architectures such as Transformer or PerceiverIO, our UAE-3D framework still achieves meaningful and comparable performances, demonstrating its robustness to different backbone choices. While further hyperparameter tuning could potentially improve the performance of these alternatives, the results already highlight the advantage of DiT in our setting.
@@
Train with RMSD ($\times 10^{-4}$Å)
@@ 6 Conclusion and Future Works
In this paper, we propose UAE-3D to compress the multi-modal features of 3D molecules into a unified latent space, and demonstrate the effectiveness of latent diffusion modeling on this space by introducing UDM-3D. By integrating atom types, chemical bonds, and 3D coordinates into a single latent space, our model effectively addresses the inherent challenges of multi-modality and SE(3) equivariance for 3D molecule generation. Extensive experiments on GEOM-Drugs and QM9 confirm leading performances in both de novo and conditional 3D molecule generation, setting new benchmarks for both quality and efficiency. UAE-3D’s success highlights the benefits of building a unified latent space for molecular design. Moving forward, we will transfer this unified latent space to new modalities, including proteins and RNAs, and extend it to broader molecular design tasks, including structure-based drug design. While UAE-3D’s latents are near-lossless compressions of the original molecule, it inspires the exploration of jointly modeling UAE-3D’s latents with text sequences for text-guided 3D molecule generation, following [45, 46, 47, 48, 49, 50, 51], as well as joint modeling of molecular and biological text representations across broader domains, including proteins [52] and single-cell transcriptomics [53].
Limitations
@@
Acknowledgements
This research/project is supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-002), the Ministry of Education (MOE T1251RES2309 and MOE T2EP20125-0039), the Agency for Science, Technology and Research (A*STAR) (H25J6a0034), and the National Natural Science Foundation of China (62572449).
Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore.
Appendix A More Experimental Results
A.1 More Ablation Studies and Analysis
@@
4 0.9201 0.8932 0.080 0.914 0.325 0.580 0.566 0.566 3.27E-01 2.20E-01 4.19E-03
8 0.9998 0.9746 0.006 0.987 0.882 0.942 0.923 0.923 1.29E-01 1.77E-02 8.14E-04
16 1.0000 1.0000 0.002 0.999 0.988 0.983 0.973 0.950 7.04E-02 9.84E-03 3.47E-04
32 1.0000 1.0000 0.003 0.986 0.867 0.943 0.918 0.918 9.98E-02 1.42E-02 7.77E-04
Table 8: The influence of R-Trans layer depth for reconstruction on the QM9 dataset.
R-Trans layers Atom Acc Bond Acc RMSD
3 1.0000 0.9999 0.00326
6 1.0000 1.0000 0.00203
9 1.0000 1.0000 0.00229
12 1.0000 1.0000 0.00209
@@
[…]
@@
Table 10: Comparison between unified and separated latent space on the QM9 dataset.
[…]
Unified 0.130 0.999 0.988 0.983 0.973 0.950 0.508 0.987 0.898
Separated 0.351 0.995 0.952 0.943 0.920 0.913 0.341 0.940 0.682
[…]
Unified 0.881 0.993 0.935 7.04E-02 9.84E-03 3.47E-04
Separated 2.356 0.982 0.872 18.4E-02 123E-03 7.03E-04
@@
Unified vs. Separated Latent Space. To validate the effectiveness of our unified latent space design, we conduct an ablation study comparing UDM-3D with a variant that employs separate latent spaces for invariant (2D) and equivariant (3D) molecular features. Specifically, we split the unified VAE into two independent neural networks: one maintaining the R-Trans architecture and taking 2D information (atom types, bond types) as input, while the other uses a vanilla transformer structure and takes 3D information (atom types and coordinates) as input. Each network generates a separate latent space, with dimensions set to half of the original hyperparameter. The LDM is then trained on the concatenated latent space of these two networks. Table 10 summarizes the results on the QM9 dataset. We observe that UDM-3D with a unified latent space significantly outperforms the separated variant across all 2D and 3D generation metrics. For instance, the FCD score improves from 0.351 to 0.130, and the MolStable metric increases from 0.952 to 0.988. This demonstrates that jointly encoding invariant and equivariant features into a single latent representation enables more effective diffusion modeling, as the LDM can learn a coherent distribution over a unified modality rather than needing to capture complex interactions between two separate modalities. These findings validate our design choice of a unified latent space as crucial for achieving state-of-the-art performance in 3D molecule generation.
Influenece of Latent Dimensions. To investigate the impact of latent dimensionality on 3D molecule generation, we conduct an ablation study on the QM9 dataset. Table 7 summarizes the reconstruction errors and the performance metrics across different latent space dimensions (4, 8, 16, and 32). The results indicate that a too-low latent dimension (e.g., 4) fails to capture the necessary information, leading to inferior performance. Increasing the dimension to 16 results in substantial performance improvements, while further increasing to 32 does not yield significant additional gains. This demonstrates that a latent dimension of 16 provides an optimal trade-off between model capacity and reconstruction fidelity.
Influence of R-Trans Layer Depth. To assess the effect of depth in the Relational Transformer encoder in UAE-3D, we conduct an ablation study by varying the number of R-Trans layers (3, 6, 9, and 12) on the QM9 dataset regarding reconstruction error. Table 8 summarizes the results. We observe that reducing the number of transformer layers to 3 leads to a noticeable increase in reconstruction error, while increasing it to 9 or 12 does not yield further improvements, as the reconstruction error is already near-zero. This indicates that a moderate depth of 6 layers is sufficient for capturing the necessary relational information among atoms and bonds for effective reconstruction.
@@
A.2 More Visualizations of Generated Molecules